OpenStudy (anonymous):

When one U.S. Dollar is worth one point seven four Australian Dollars, which of the following could be used as a conversion factor? five hundred eighty-sevenths one and sixteen twenty-fifths one and thirty-seven fiftieths two and thirteen fiftieths

OpenStudy (austinl):

Try converting 1.74 to a mixed number?

OpenStudy (anonymous):

im not that smart D:

OpenStudy (austinl):

\[1.74 = \frac{174}{100}\] \[\frac{174}{2} = ?\] \[\frac{100}{2} = ?\]

OpenStudy (anonymous):

good explaining @austinL

OpenStudy (austinl):

\[\frac{174}{2} = x\] \[\frac{100}{2} = y\] \[1.74 = \frac{x}{y}\] Then you turn x/y into a mixed number.

OpenStudy (anonymous):

thankss

OpenStudy (austinl):

No problem :)